# Fairness-Aware Domain Generalization under Covariate and Dependence Shifts

### v1

@@ C.4 Learning the Transformation Model
Adversarial Loss. Motivated by the observation that GANs [46] can improve data quality for evaluating the disentanglement effect in the latent spaces, we use GANs to match the distribution of reconstructed data to the same distribution. Followed by [18], data and semantic factors generated through encoders and decoders should be indistinguishable from the given ones in the same domain.
$\mathcal{L}_{adv}$ $=\underbrace{\mathbb{E}_{\mathbf{c}\sim p(\mathbf{c}),\mathbf{s}\sim\mathcal{N}(0,\mathbf{I}_{s}),\mathbf{a}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{a})}\big{[}\log\big{(}1-D^{o}(G^{o}(\hat{\mathbf{m}},\mathbf{s}))\big{)}\big{]}+\mathbb{E}_{\mathbf{x}\sim p(\mathbf{x})}\big{[}\log D^{o}(\mathbf{x})\big{]}}_{\text{outer level}}$
$+\underbrace{\mathbb{E}_{\mathbf{c}\sim p(\mathbf{c}),\mathbf{a}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{a})}\big{[}\log\big{(}1-D^{i}(G^{i}(\mathbf{c,a}))\big{)}\big{]}+\mathbb{E}_{\mathbf{m}\sim p(\mathbf{m})}\big{[}\log D^{i}(\mathbf{m})\big{]}}_{\text{inner level}}$
@@
where $D^{o}:\mathcal{X}\times\Psi\rightarrow\mathbb{R}$ and $D^{i}:\mathcal{M}\times\Psi\rightarrow\mathbb{R}$ are the discriminators for the outer and inner levels parameterized by $\bm{\psi}_{o}\in\Psi$ and $\bm{\psi}_{i}\in\Psi$, respectively.
@@
$\min_{E^{m},E^{s},E^{c},E^{a},G^{i},G^{o}}\>\max_{D^{i},D^{o}}\>\mathcal{L}_{total}=\beta_{d}\mathcal{L}^{data}_{recon}+\beta_{f}\mathcal{L}^{factor}_{recon}+\beta_{z}\mathcal{L}_{sens}+\beta_{g}\mathcal{L}_{adv}$ (9)
where $\beta_{d},\beta_{f},\beta_{z},\beta_{g}>0$ are hyperparameters that control the importance of each loss term. To optimize, the learning algorithm is given in Algorithm 2.
@@
Require: learning rate $\alpha_{1},\alpha_{2},\alpha_{3}$, initial coefficients $\beta_{d},\beta_{f},\beta_{z},\beta_{g}$.
Initialize: Parameter of encoders $\{\bm{\theta}_{m},\bm{\theta}_{s},\bm{\theta}_{c},\bm{\theta}_{a}$}, decoders $\{\bm{\phi}_{i},\bm{\phi}_{o}\}$, sensitive classifier $\bm{\theta}_{z}$, and discriminators $\{\bm{\psi}_{i},\bm{\psi}_{o}\}$.
@@
4:         $\bm{\psi}_{o},\bm{\psi}_{i}\leftarrow\text{Adam}(\beta_{g}\mathcal{L}_{adv},\bm{\psi}_{o},\bm{\psi}_{i},\alpha_{1})$
@@
$\bm{\theta}_{m},\bm{\theta}_{c},\bm{\theta}_{s},\bm{\theta}_{a},\bm{\phi}_{o},\bm{\phi}_{i}\leftarrow\text{Adam}\big{(}\beta_{d}\mathcal{L}^{data}_{recon}+\beta_{f}\mathcal{L}^{factor}_{recon},\bm{\theta}_{m},\bm{\theta}_{c},\bm{\theta}_{s},\bm{\theta}_{a},\bm{\phi}_{o},\bm{\phi}_{i},\alpha_{2}\big{)}$
6:         $\bm{\theta}_{z}\leftarrow\text{Adam}(\beta_{z}\mathcal{L}_{sens},\bm{\theta}_{z},\alpha_{3})$
@@
9:Return $\{\bm{\theta}_{m},\bm{\theta}_{s},\bm{\theta}_{c},\bm{\theta}_{a},\bm{\theta}_{z},\bm{\phi}_{i},\bm{\phi}_{o}\}$
C.5 Architecture Details
@@
For ccMNIST, FairFace, and YFCC100M-FDG datasets: All the images are resized to $224\times 224$. $E^{m}$ and $E^{c}$’s structures are the same. Each of them is made of four convolution layers. The first one has $64$ filters, and each of the others has $128$ filters. The kernel sizes are $(7,7),(4,4),(3,3),(3,3)$ for layers $1$ to $4$, respectively. The stride of the second layer is $(2,2)$, and the stride of all the other layers is $(1,1)$. The activation function of the first three layers is ReLU. The last convolution layer does not have an activation function. $E^{s}$ and $E^{a}$’s structures are the same. Each of them is made of $6$ convolution layers, and there is an adaptive average pooling layer with output size $1$ between the last two convolution layers. The numbers of filters are $64,128,256,256,256$, and $2$ for the convolution layers, respectively. The kernel sizes are $(7,7),(4,4),(4,4),(4,4),(4,4),(1,1)$. And the strides are $(1,1),(2,2),(2,2),(2,2),(2,2),(1,1)$. The activation function of the first five layers is ReLU. The last convolution layer does not have an activation function. $G^{o}$ and $G^{i}$’s structures are almost the same. The only difference between them is the output size, $3$ for $G^{o}$ and $128$ for $G^{i}$. Each of them has two parts. The first part is made of $4$ convolution layers, and there is an upsampling layer with a scale factor $2.0$ between the second convolution layer and the third convolution layer. The numbers of filters are $128,128,64,$ and $3$ for the convolution layers, respectively. The kernel sizes are $(3,3),(3,3),(5,5),(7,7)$. The strides are $(1,1)$ for all the convolution layers. The first and the third convolution layers’ activation functions are ReLU. The fourth convolution layer’s activation function is Tanh. The second convolution layer does not have an activation function. The second part is made of three fully connected layers. The number of neurons is $256$ and $256$, respectively, and the output size is $512$. The activation function of the first two layers is ReLU, and there is no activation function on the output. $D^{o}$ comprises $4$ convolution layers followed by an average pooling layer whose kernel size is $3$, stride is $2$, and padding is $[1,1]$. The numbers of filters of the convolution layers are $64,128,256,1$, respectively. The kernel sizes are $(4,4)$ for the first three convolution layers and $(1,1)$ for the fourth convolution layer. The strides are $(2,2)$ for the first three convolution layers and $(1,1)$ for the fourth convolution layer. The first three convolution layers’ activation functions are LeakyReLU. The other layers do not have activation functions. $D^{i}$ is made of one fully connected layer whose input size is $112$, and the output size is $64$ with activation function ReLU. $h$ comprises one fully connected layer with input size $2$, output size $1$, and activation function Sigmoid. $f$ has two parts. The first part is Resnet-50 [47], and the second is one fully connected layer with input size $2048$ and output size $2$.
@@
For the NYSF dataset: $E^{m}$ is made of two fully connected layers. The number of neurons is $32$, and the output size is $16$. The activation function of the first layer is ReLU, and there is no activation function on the output. $E^{s}$ is made of two fully connected layers. The number of neurons is $32$, and the output size is $2$. The activation function of the first layer is ReLU, and there is no activation function on the output. $G^{o}$ is made of two fully connected layers. The number of neurons is $32$, and the output size is $51$. The activation function of the first layer is ReLU, and there is no activation function on the output. $D^{o}$ is made of two fully connected layers. The number of neurons is $32$, and the output size is $16$. The activation function of the first layer is ReLU, and there is no activation function on the output. $E^{c}$ is made of two fully connected layers. The number of neurons is $16$, and the output size is $8$. The activation function of the first layer is ReLU, and there is no activation function on the output. $E^{a}$ is made of two fully connected layers. The number of neurons is $8$, and the output size is $2$. The activation function of the first layer is ReLU, and there is no activation function on the output. $G^{i}$ is made of two fully connected layers. The number of neurons is $16$, and the output size is $16$. The activation function of the first layer is ReLU, and there is no activation function on the output. $D^{i}$ is made of two fully connected layers. The number of neurons is $8$, and the output size is $8$. The activation function of the first layer is ReLU, and there is no activation function on the output. $h$ comprises one fully connected layer with input size $2$ and output size $1$. The activation function is Sigmoid. $f$ has two parts. The first part is made of $3$ fully connected layers. The number of neurons is $32$, and the output size is $32$. The activation function of the first two layers is ReLU, and there is no activation function on the output. The second part is made of one fully connected layer whose input size is $32$, the output size is $32$, and it does not have an activation function.
C.6 Hyperparameter Search
@@
We follow the same set of the MUNIT [18] for the hyperparameters. More specifically, the learning rate is $0.0001$, the number of iterations is $600000$, and the batch size is $1$. The loss weights in learning $T$ are chosen from $\{1,5,10\}$. The selected best ones are $\beta_{d}=10,\beta_{f}=1,\beta_{z}=1,\beta_{g}=1$. We monitor the loss of the validation set and choose the $\beta$ with the lowest validation loss.
@@
For the hyperparameters in learning the classifier $f$, the learning rate is chosen from $\{0.000005,0.00001,0.00005,0.0001,0.0005\}$. $\eta$ is chosen from $\{0.01,0.05,0.1\}$. $\gamma$ is chosen from $\{0.01,0.025,0.05\}$. $\lambda$ is chosen from $\{0.1,1,10,20\}$. The batch size is chosen from $\{22,64,80,128,512,1024,2048\}$. The numbers of iterations are chosen from $\{500,1000,...,8000\}$ on the ccMNIST and NYSF datasets. The number of iterations are chosen from $\{300,600,...,7800,8000\}$ on the FairFace and YFCC100M-FDG datasets. The selected best ones are: the learning rate is $0.00005$, $\eta_{1}=\eta_{2}=0.05$, $\gamma_{1}=\gamma_{2}=0.025$, $\lambda_{1}=\lambda_{2}=1$. The batch size on the ccMNIST and YFCC100M-FDG datasets is $64$, and it is $22$ on the FairFace dataset and $1024$ on the NYSF dataset. The number of iterations on the ccMNIST dataset is $3000,500,7000$ for domains R, G, B, respectively. The number of iterations on the FairFace dataset is $7200,7200,7800,8000,6600,7200,6900$ for domains B, E, I, L, M, S, W, respectively. The number of iterations on the YFCC100M-FDG dataset is $7200,6000,6900$ for $d_{0},d_{1},d_{2}$, respectively. The number of iterations on the NYSF dataset is $500,3500,4000,1500,8000$ for domains R, B, M, Q, S, respectively. We monitor the accuracy and the value of fairness metrics from the validation set and select the best ones. The grid space of the grid search on all the baselines is the same as for our method.
C.7 Model Selection.
The model selection in domain generalization is intrinsically a learning problem, followed by [12], we use leave-one-domain-out validation criteria, which is one of the three selection methods stated in [48]. Specifically, we evaluate FDDG on the held-out training domain and average the performance of $|\mathcal{E}_{s}|-1$ domains over the held-out one.
Appendix D Ablation Studies
@@
We conduct three ablation studies, and detailed algorithms of designed ablation studies are given in Algorithms 3, 4 and 5. For additional ablation study results on ccMNIST, YFCC100M-FDG, and NYSF, refer to Appendix F.
@@
    The difference between the full FDDG and the first ablation study (FDDG w/o sf) is that the latter does not have the inner level when learning $T$. Since the inner level is used to extract the content and sensitive factors from the semantic one, the same sensitive label of the generated images will remain due to the absence of $h(\cdot)$. Therefore, FDDG w/o sf is expected to have a lower level of fairness in the experiments. Results shown in the tables indicate that FDDG w/o sf has a significantly lower performance on fairness metrics.
@@
    The second study (FDDG w/o $T$) does not train the auto-encoders to generate images. All losses are computed only based on the sampled images. Similar to FDDG w/o sf, it is much harder to train a good classifier without the generated images in synthetic domains. Our results demonstrate that FDDG w/o $T$ performs worse on all the datasets.
@@
    The difference between FDDG and the third study (FDDG w/o fc) is that FDDG w/o fc does not have the fairness loss $\mathcal{L}_{fair}$ in line 9 of Algorithm 1. Therefore, this algorithm only focuses on accuracy without considering fairness. Results based on FDDG w/o fc show that it has a good level of accuracy but a poor level of fairness.
Algorithm 3 FDDG w/o sf (Ablation Study 1)
@@
3:         $\mathcal{L}_{cls}(\bm{\theta})=(1/m)\sum_{i=1}^{m}\ell(y_{i},\hat{f}(\mathbf{x}_{i},\bm{\theta}))$
@@
$\mathcal{L}_{fair}(\bm{\theta})=(1/m)\sum_{i=1}^{m}(\frac{1}{p_{1}(1-p_{1})}(\frac{z_{i}+1}{2}-p_{1})\hat{f}(\mathbf{x}_{i},\bm{\theta})$
@@
$(\mathbf{x}^{\prime}_{i},z_{i},y_{i})=\textsc{DataAug}(\mathbf{x}_{i},x_{i},y_{i})$
7:              $\mathcal{L}^{\prime}_{cls}(\bm{\theta})=(1/m)\sum_{i=1}^{m}\ell(y_{i},\hat{f}(\mathbf{x}^{\prime}_{i},\bm{\theta}))$
@@
$\mathcal{L}_{cls}(\bm{\theta})=\mathcal{L}_{cls}(\bm{\theta})+\mathcal{L}^{\prime}_{cls}(\bm{\theta})$
10:         $\mathcal{L}(\bm{\theta})=\mathcal{L}_{cls}(\bm{\theta})+\lambda_{2}\cdot\mathcal{L}_{fair}(\bm{\theta})$
11:         $\bm{\theta}\leftarrow\bm{\theta}-\eta_{p}\cdot\nabla_{\bm{\theta}}\mathcal{L}(\bm{\theta})$
12:         $\lambda_{2}\leftarrow\text{max}\{[\lambda_{2}+\eta_{d}\cdot(\mathcal{L}_{fair}(\bm{\theta})-\gamma_{2})],0\}$
@@
15:procedure DataAug($\mathbf{x},z,y$)
16:     $\mathbf{c}=E^{m}(\mathbf{x},\bm{\theta}^{m})$
@@
18:     $\mathbf{x}^{\prime}=G^{o}(\mathbf{c},\mathbf{s}^{\prime},\bm{\phi}_{o})$
@@
Algorithm 4 FDDG w/o $T$ (Ablation Study 2)
@@
3:         $\mathcal{L}_{cls}(\bm{\theta})=(1/m)\sum_{i=1}^{m}\ell(y_{i},\hat{f}(\mathbf{x}_{i},\bm{\theta}))$
@@
$\mathcal{L}_{fair}(\bm{\theta})=(1/m)\sum_{i=1}^{m}(\frac{1}{p_{1}(1-p_{1})}(\frac{z_{i}+1}{2}-p_{1})\hat{f}(\mathbf{x}_{i},\bm{\theta})$
5:         $\mathcal{L}(\bm{\theta})=\mathcal{L}_{cls}(\bm{\theta})+\lambda_{2}\cdot\mathcal{L}_{fair}(\bm{\theta})$
6:         $\bm{\theta}\leftarrow\bm{\theta}-\eta_{p}\cdot\nabla_{\bm{\theta}}\mathcal{L}(\bm{\theta})$
7:         $\lambda_{2}\leftarrow\text{max}\{[\lambda_{2}+\eta_{d}\cdot(\mathcal{L}_{fair}(\bm{\theta})-\gamma_{2})],0\}$
@@
Algorithm 5 FDDG w/o fc (Ablation Study 3)
@@
3:         $\mathcal{L}_{cls}(\bm{\theta})=(1/m)\sum_{i=1}^{m}\ell(y_{i},\hat{f}(\mathbf{x}_{i},\bm{\theta}))$
4:         Initialize $\mathcal{L}^{\prime}_{inv}(\bm{\theta})=0$
@@
$(\mathbf{x}^{\prime}_{i},y_{i})=\textsc{DataAug}(\mathbf{x}_{i},z_{i},y_{i})$
@@
$\mathcal{L}^{\prime}_{inv}(\bm{\theta})\mathrel{+}=d[\hat{f}(\mathbf{x}_{i},\bm{\theta}),\hat{f}(\mathbf{x}^{\prime}_{i},\bm{\theta})]$

### v2

@@ B.2 Details of Learning the Transformation Model
Additionally, motivated by the observation that GANs [65] can improve data quality for evaluating the disentanglement effect in the latent spaces, we use GANs to match the distribution of reconstructed data to the same distribution. Followed by [48], data and semantic factors generated through encoders and decoders should be indistinguishable from the given ones in the same domain.
$\mathcal{L}_{adv}=$ $\mathbb{E}_{\mathbf{c}\sim\mathbb{P}_{C},\mathbf{s}^{s}\sim\mathcal{N}(0,\mathbf{I}_{s}),\mathbf{a}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{a})}\big{[}\log\big{(}1-D_{o}(G_{o}(\hat{\mathbf{m}},\mathbf{s}^{s}))\big{)}\big{]}+\mathbb{E}_{\mathbf{x}^{s}\sim\mathbb{P}^{s}_{X}}\big{[}\log D_{o}(\mathbf{x}^{s})\big{]}$
$+\mathbb{E}_{\mathbf{c}\sim\mathbb{P}_{C},\mathbf{a}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{a})}\big{[}\log\big{(}1-D_{i}(G_{i}(\mathbf{c,a}))\big{)}\big{]}+\mathbb{E}_{\mathbf{m}\sim\mathbb{P}_{M}}\big{[}\log D_{i}(\mathbf{m})\big{]}$
@@
where $D_{o}:\mathcal{X}\rightarrow\mathbb{R}$ and $D_{i}:\mathcal{M}\rightarrow\mathbb{R}$ are the discriminators for the outer and inner levels, respectively.
@@
$\min_{E_{m},E_{s},E_{c},E_{a},G_{i},G_{o}}\>\max_{D_{i},D_{o}}\>\beta_{1}\mathcal{L}^{data}_{recon}+\beta_{2}\mathcal{L}^{factor}_{recon}+\beta_{3}\mathcal{L}_{sens}+\beta_{4}\mathcal{L}_{adv}$ (10)
where $\beta_{1},\beta_{2},\beta_{3},\beta_{4}>0$ are hyperparameters that control the importance of each loss term. To optimize, the learning algorithm is given in Algorithm 2.
@@
Require: learning rate $\alpha_{1},\alpha_{2},\alpha_{3}$, initial coefficients $\beta_{1},\beta_{2},\beta_{3},\beta_{4}$.
Initialize: Parameter of encoders $\{\boldsymbol{\theta}_{m},\boldsymbol{\theta}_{s},\boldsymbol{\theta}_{c},\boldsymbol{\theta}_{a}$}, decoders $\{\boldsymbol{\phi}_{i},\boldsymbol{\phi}_{o}\}$, sensitive classifier $\boldsymbol{\theta}_{z}$, and discriminators $\{\boldsymbol{\psi}_{i},\boldsymbol{\psi}_{o}\}$.
@@
4:         $\boldsymbol{\psi}_{o},\boldsymbol{\psi}_{i}\leftarrow\text{Adam}(\beta_{4}\mathcal{L}_{adv},\boldsymbol{\psi}_{o},\boldsymbol{\psi}_{i},\alpha_{1})$
@@
$\boldsymbol{\theta}_{m},\boldsymbol{\theta}_{c},\boldsymbol{\theta}_{s},\boldsymbol{\theta}_{a},\boldsymbol{\phi}_{o},\boldsymbol{\phi}_{i}\leftarrow\text{Adam}\big{(}\beta_{1}\mathcal{L}^{data}_{recon}+\beta_{2}\mathcal{L}^{factor}_{recon},\boldsymbol{\theta}_{m},\boldsymbol{\theta}_{c},\boldsymbol{\theta}_{s},\boldsymbol{\theta}_{a},\boldsymbol{\phi}_{o},\boldsymbol{\phi}_{i},\alpha_{2}\big{)}$
6:         $\boldsymbol{\theta}_{z}\leftarrow\text{Adam}(\beta_{3}\mathcal{L}_{sens},\boldsymbol{\theta}_{z},\alpha_{3})$
@@
9:Return $\{\boldsymbol{\theta}_{m},\boldsymbol{\theta}_{s},\boldsymbol{\theta}_{c},\boldsymbol{\theta}_{a},\boldsymbol{\theta}_{z},\boldsymbol{\phi}_{i},\boldsymbol{\phi}_{o}\}$
B.3 Architecture Details
@@
For ccMNIST, FairFace, and YFCC100M-FDG datasets: All the images are resized to $224\times 224$. $E_{m}$ and $E_{c}$’s structures are the same. Each of them is made of four convolution layers. The first one has $64$ filters, and each of the others has $128$ filters. The kernel sizes are $(7,7),(4,4),(3,3),(3,3)$ for layers $1$ to $4$, respectively. The stride of the second layer is $(2,2)$, and the stride of all the other layers is $(1,1)$. The activation function of the first three layers is ReLU. The last convolution layer does not have an activation function. $E_{s}$ and $E_{a}$’s structures are the same. Each of them is made of $6$ convolution layers, and there is an adaptive average pooling layer with output size $1$ between the last two convolution layers. The numbers of filters are $64,128,256,256,256$, and $2$ for the convolution layers, respectively. The kernel sizes are $(7,7),(4,4),(4,4),(4,4),(4,4),(1,1)$. And the strides are $(1,1),(2,2),(2,2),(2,2),(2,2),(1,1)$. The activation function of the first five layers is ReLU. The last convolution layer does not have an activation function. $G_{o}$ and $G_{i}$’s structures are almost the same. The only difference between them is the output size, $3$ for $G_{o}$ and $128$ for $G_{i}$. Each of them has two parts. The first part is made of $4$ convolution layers, and there is an upsampling layer with a scale factor $2.0$ between the second convolution layer and the third convolution layer. The numbers of filters are $128,128,64,$ and $3$ for the convolution layers, respectively. The kernel sizes are $(3,3),(3,3),(5,5),(7,7)$. The strides are $(1,1)$ for all the convolution layers. The first and the third convolution layers’ activation functions are ReLU. The fourth convolution layer’s activation function is Tanh. The second convolution layer does not have an activation function. The second part is made of three fully connected layers. The number of neurons is $256$ and $256$, respectively, and the output size is $512$. The activation function of the first two layers is ReLU, and there is no activation function on the output. $D_{o}$ comprises $4$ convolution layers followed by an average pooling layer whose kernel size is $3$, stride is $2$, and padding is $[1,1]$. The numbers of filters of the convolution layers are $64,128,256,1$, respectively. The kernel sizes are $(4,4)$ for the first three convolution layers and $(1,1)$ for the fourth convolution layer. The strides are $(2,2)$ for the first three convolution layers and $(1,1)$ for the fourth convolution layer. The first three convolution layers’ activation functions are LeakyReLU. The other layers do not have activation functions. $D_{i}$ is made of one fully connected layer whose input size is $112$, and the output size is $64$ with the activation function ReLU. $h$ comprises one fully connected layer with input size $2$, output size $1$, and activation function Sigmoid. $f$ has two parts. The first part is Resnet-50 [66], and the second is one fully connected layer with input size $2048$ and output size $2$.
@@
For the NYSF dataset: $E_{m}$ is made of two fully connected layers. The number of neurons is $32$, and the output size is $16$. The activation function of the first layer is ReLU, and there is no activation function on the output. $E_{s}$ is made of two fully connected layers. The number of neurons is $32$, and the output size is $2$. The activation function of the first layer is ReLU, and there is no activation function on the output. $G_{o}$ is made of two fully connected layers. The number of neurons is $32$, and the output size is $51$. The activation function of the first layer is ReLU, and there is no activation function on the output. $D_{o}$ is made of two fully connected layers. The number of neurons is $32$, and the output size is $16$. The activation function of the first layer is ReLU, and there is no activation function on the output. $E_{c}$ is made of two fully connected layers. The number of neurons is $16$, and the output size is $8$. The activation function of the first layer is ReLU, and there is no activation function on the output. $E_{a}$ is made of two fully connected layers. The number of neurons is $8$, and the output size is $2$. The activation function of the first layer is ReLU, and there is no activation function on the output. $G_{i}$ is made of two fully connected layers. The number of neurons is $16$, and the output size is $16$. The activation function of the first layer is ReLU, and there is no activation function on the output. $D_{i}$ is made of two fully connected layers. The number of neurons is $8$, and the output size is $8$. The activation function of the first layer is ReLU, and there is no activation function on the output. $h$ comprises one fully connected layer with input size $2$ and output size $1$. The activation function is Sigmoid. $f$ has two parts. The first part is made of $3$ fully connected layers. The number of neurons is $32$, and the output size is $32$. The activation function of the first two layers is ReLU, and there is no activation function on the output. The second part is made of one fully connected layer whose input size is $32$, the output size is $32$, and it does not have an activation function.
B.4 Hyperparameter Search
@@
We follow the same set of the MUNIT [48] for the hyperparameters. More specifically, the learning rate is $0.0001$, the number of iterations is $600000$, and the batch size is $1$. The loss weights in learning $T$ are chosen from $\{1,5,10\}$. The selected best ones are $\beta_{1}=10,\beta_{2}=1,\beta_{3}=1,\beta_{4}=1$. We monitor the loss of the validation set and choose the $\beta$ with the lowest validation loss.
@@
For the hyperparameters in learning the classifier $f$, the learning rate is chosen from $\{0.000005,0.00001,0.00005,0.0001,0.0005\}$. $\eta$ is chosen from $\{0.01,0.05,0.1\}$. $\gamma$ is chosen from $\{0.01,0.025,0.05\}$. $\lambda$ is chosen from $\{0.1,1,10,20\}$. The batch size is chosen from $\{22,64,80,128,512,1024,2048\}$. The numbers of iterations is chosen from $\{500,1000,...,8000\}$ on the ccMNIST and NYSF datasets. The number of iterations are chosen from $\{300,600,...,7800,8000\}$ on the FairFace and YFCC100M-FDG datasets. The selected best ones are: the learning rate is $0.00005$, $\eta_{1}=\eta_{2}=0.05$, $\gamma_{1}=\gamma_{2}=0.025$, $\lambda_{1}=\lambda_{2}=1$. The batch size on the ccMNIST and YFCC100M-FDG datasets is $64$, and it is $22$ on the FairFace dataset and $1024$ on the NYSF dataset. The number of iterations on the ccMNIST dataset is $3000,500,7000$ for domains R, G, B, respectively. The number of iterations on the FairFace dataset is $7200,7200,7800,8000,6600,7200,6900$ for domains B, E, I, L, M, S, W, respectively. The number of iterations on the YFCC100M-FDG dataset is $7200,6000,6900$ for $d_{0},d_{1},d_{2}$, respectively. The number of iterations on the NYSF dataset is $500,3500,4000,1500,8000$ for domains R, B, M, Q, S, respectively. We monitor the accuracy and the value of fairness metrics from the validation set and select the best ones. The grid space of the grid search on all the baselines is the same as for our method.
Appendix C Ablation Studies
@@
We conduct three ablation studies, and detailed algorithms of designed ablation studies are given in Algorithms 3, 4 and 5. For additional ablation study results on ccMNIST, YFCC100M-FDG, and NYSF, refer to Appendix E.
@@
    The difference between the full FEDORA and the first ablation study (w/o $E_{a}$) is that the latter does not have the inner level when learning $T$. Since the inner level is used to extract the content and sensitive factors from the semantic one, the same sensitive label of the generated images will remain due to the absence of $h(\cdot)$. Therefore, w/o $E_{a}$ is expected to have a lower level of fairness in the experiments. Results shown in the tables indicate that w/o $E_{a}$ has a significantly lower performance on fairness metrics.
@@
    The second study (w/o $T$) does not train the auto-encoders to generate images. All losses are computed only based on the sampled images. Similar to w/o $E_{a}$, it is much harder to train a good classifier without the generated images in synthetic domains. Our results demonstrate that w/o $T$ performs worse on all the datasets.
@@
    The difference between FEDORA and the third study (w/o $\mathcal{L}_{fair}$) is that w/o $\mathcal{L}_{fair}$ does not have the fairness loss $\mathcal{L}_{fair}$ in line 9 of Algorithm 1. Therefore, this algorithm only focuses on accuracy without considering fairness. Results based on w/o $\mathcal{L}_{fair}$ show that it has a good level of accuracy but a poor level of fairness.
Algorithm 3 w/o $E_{a}$ (Ablation Study 1)
@@
3:         $\mathcal{L}_{cls}(\boldsymbol{\theta})=(1/m)\sum_{i=1}^{m}\ell(y_{i},\hat{f}(\mathbf{x}_{i},\boldsymbol{\theta}))$
@@
$\mathcal{L}_{fair}(\boldsymbol{\theta})=(1/m)\sum_{i=1}^{m}(\frac{1}{p_{1}(1-p_{1})}(\frac{z_{i}+1}{2}-p_{1})\hat{f}(\mathbf{x}_{i},\boldsymbol{\theta})$
@@
$(\mathbf{x}^{\prime}_{i},z_{i},y_{i})=\textsc{T}(\mathbf{x}_{i},x_{i},y_{i})$
7:              $\mathcal{L}^{\prime}_{cls}(\boldsymbol{\theta})=(1/m)\sum_{i=1}^{m}\ell(y_{i},\hat{f}(\mathbf{x}^{\prime}_{i},\boldsymbol{\theta}))$
@@
$\mathcal{L}_{cls}(\boldsymbol{\theta})=\mathcal{L}_{cls}(\boldsymbol{\theta})+\mathcal{L}^{\prime}_{cls}(\boldsymbol{\theta})$
10:         $\mathcal{L}(\boldsymbol{\theta})=\mathcal{L}_{cls}(\boldsymbol{\theta})+\lambda_{2}\cdot\mathcal{L}_{fair}(\boldsymbol{\theta})$
11:         $\boldsymbol{\theta}\leftarrow\boldsymbol{\theta}-\eta_{p}\cdot\nabla_{\boldsymbol{\theta}}\mathcal{L}(\boldsymbol{\theta})$
12:         $\lambda_{2}\leftarrow\text{max}\{[\lambda_{2}+\eta_{d}\cdot(\mathcal{L}_{fair}(\boldsymbol{\theta})-\gamma_{2})],0\}$
@@
15:procedure T($\mathbf{x},z,y$)
16:     $\mathbf{c}=E^{m}(\mathbf{x},\boldsymbol{\theta}^{m})$
@@
18:     $\mathbf{x}^{\prime}=G^{o}(\mathbf{c},\mathbf{s}^{\prime},\boldsymbol{\phi}_{o})$
@@
Algorithm 4 w/o $T$ (Ablation Study 2)
@@
3:         $\mathcal{L}_{cls}(\boldsymbol{\theta})=(1/m)\sum_{i=1}^{m}\ell(y_{i},\hat{f}(\mathbf{x}_{i},\boldsymbol{\theta}))$
@@
$\mathcal{L}_{fair}(\boldsymbol{\theta})=(1/m)\sum_{i=1}^{m}(\frac{1}{p_{1}(1-p_{1})}(\frac{z_{i}+1}{2}-p_{1})\hat{f}(\mathbf{x}_{i},\boldsymbol{\theta})$
5:         $\mathcal{L}(\boldsymbol{\theta})=\mathcal{L}_{cls}(\boldsymbol{\theta})+\lambda_{2}\cdot\mathcal{L}_{fair}(\boldsymbol{\theta})$
6:         $\boldsymbol{\theta}\leftarrow\boldsymbol{\theta}-\eta_{p}\cdot\nabla_{\boldsymbol{\theta}}\mathcal{L}(\boldsymbol{\theta})$
7:         $\lambda_{2}\leftarrow\text{max}\{[\lambda_{2}+\eta_{d}\cdot(\mathcal{L}_{fair}(\boldsymbol{\theta})-\gamma_{2})],0\}$
@@
Algorithm 5 w/o $\mathcal{L}_{fair}$ (Ablation Study 3)
@@
3:         $\mathcal{L}_{cls}(\boldsymbol{\theta})=(1/m)\sum_{i=1}^{m}\ell(y_{i},\hat{f}(\mathbf{x}_{i},\boldsymbol{\theta}))$
4:         Initialize $\mathcal{L}^{\prime}_{inv}(\boldsymbol{\theta})=0$
@@
$(\mathbf{x}^{\prime}_{i},y_{i})=\textsc{T}(\mathbf{x}_{i},z_{i},y_{i})$
@@
$\mathcal{L}^{\prime}_{inv}(\boldsymbol{\theta})\mathrel{+}=d[\hat{f}(\mathbf{x}_{i},\boldsymbol{\theta}),\hat{f}(\mathbf{x}^{\prime}_{i},\boldsymbol{\theta})]$